A presidential candidate asks a polling organization to conduct a nationwide survey to determine the percentage of potential voters who would vote for him over his rival presidential candidate. Out of 2500 respondents in the sample, 925 said they would vote for him. If 40% of the potential voters vote for his rival, is this significantly different from the percentage of potential voters of the candidate who requested the survey? Use 5% level of significance.


1
Expert's answer
2021-05-07T09:54:50-0400

n = 2500

z = 1.960

p = 925 / 2500 = 0.37

0.37 "\pm1.96\sqrt{\frac{0.37\times0.63}{2500}}" = 0.37"\pm" 0.0189

Upper limit is 0.37 + 0.0189 = 0.3889

Answer: at 0.05 level of significance there is a difference.